Find the form of expression of the period of vibration of simple pendulum as T=km^x L^y G^2 where k,x,G, and 2 are known terms?

1
Expert's answer
2021-11-03T10:17:12-0400

answer:-

"T=km^xl^yg^z"

"[MLT]=Km^xL^yG^z\\
g=m/s^2\\"

where k is dimensionless constant of proportionality.

Writing the dimensions in terms of M, L, T on either side , we get

"[M^0L^0T^1]=K[M]^x[L]^{y+z}[T]^z"

Applying the principle of homogeneity of dimensions, we get

"x=0\\
y=\frac{1}{2}\\
z=\frac{-1}{2}"

so ,

"T=K \sqrt{\frac{L}{G}}"